# Compatible $E$-differential forms on Lie algebroids over (pre-)multisymplectic manifolds

### v3

@@ 7.2 QP-manifold description of compatible differential forms
Our graded manifold is $T^{*}[n]{\mathcal{M}}=T^{*}[n]T^{*}[n-1]E[1]$. In order to compute the concrete formula, take local coordinates $(x^{i},q^{a},z_{i},y_{a})$ of degree $(0,1,n-1,n-2)$ on $T^{*}[n-1]E[1]$ and canonical conjugate coordinates of the fiber $(\xi_{i},p_{a},\zeta^{i},\eta^{a})$ of degree $(n,n-1,1,2)$. The canonical projection ${pr}:T^{*}[n]{\mathcal{M}}\rightarrow{\mathcal{M}}$ is given by ${pr}:(\xi_{i},p_{a},\zeta^{i},\eta_{a})\mapsto(0,0,0,0)$. The canonical symplectic form of degree $n$ is
${\widetilde{\omega}}_{grad}$ $=\delta x^{i}\wedge\delta\xi_{i}+\delta q^{a}\wedge\delta p_{a}+\delta z_{i}\wedge\delta\zeta^{i}+\delta y_{a}\wedge\delta\eta^{a}.$ (93)
@@
$\{x^{i},\xi_{j}\}$ =$=-\{\xi_{j},x^{i}\}=\delta^{i}_{j},$ (94)
$\{q^{a},p_{b}\}$ $=-(-1)^{n-1}\{p_{b},q^{a}\}=\delta^{a}_{b},$ (95)
$\{z_{i},\zeta^{j}\}$ $=-(-1)^{n-1}\{\zeta^{j},z_{i}\}=\delta_{i}^{j},$ (96)
$\{y_{a},\eta^{b}\}$ =$=-\{\eta^{b},y_{a}\}=\delta_{a}^{b}.$ (97)
We take an $n+1$-form $H$ on $M$ and consider the following function of degree $n+1$,
@@
$=\xi_{i}\zeta^{i}+p_{a}\eta^{a}+(-1)^{\tfrac{n(n-1)}{2}}\frac{1}{(n+1)!}H_{i_{1}\ldots i_{n+1}}(x)\zeta^{i_{1}}\ldots\zeta^{i_{n+1}},$ (98)
@@
where $\tilde{H}=\frac{1}{(n+1)!}H_{i_{1}\ldots i_{n+1}}(x)\zeta^{i_{1}}\ldots\zeta^{i_{n+1}}$ is a degree $n+1$ function corresponding to the $n+1$-form $H=\frac{1}{(n+1)!}H_{i_{1}\ldots i_{n+1}}(x)\mathrm{d}x^{i_{1}}\wedge\ldots\wedge\mathrm{d}x^{i_{n+1}}\in\Omega^{n+1}(M)$. An $n+1$-form is identified to a degree $n+1$ function on $T[1]M$. ${\widetilde{\Theta}}$ satisfies ${\{{{{\widetilde{\Theta}}},{{\widetilde{\Theta}}}}\}}=0$ if and only if $H$ is closed, $\mathrm{d}H=0$, i.e., $(T^{*}[n]{\mathcal{M}},{\widetilde{\omega}}_{grad},{\widetilde{Q}})$ is a QP-manifold if $H$ is a pre-$n$-plectic form on $M$, where ${\widetilde{Q}}$ is a Hamiltonian vector field for ${\widetilde{\Theta}}$.
@@
We consider a twist using the following function of degree $n$, $\phi={\phi_{Lie}}-{\underline{J}}{}$, where two functions are
${\phi_{Lie}}$ $=\rho^{i}_{a}(x)z_{i}q^{a}+\frac{1}{2}C^{c}_{ab}(x)q^{a}q^{b}y_{c},$ (99)
${\underline{J}}{}$ $=\frac{1}{n!}J_{a_{1}\ldots a_{n}}(x)q^{a_{1}}\ldots q^{a_{n}}.$ (100)
@@
and impose the condition (92), ${pr}_{*}e^{\mathrm{ad}\phi}{\widetilde{\Theta}}=0$. ${\phi_{Lie}}$ is a similar form as $\Theta_{Lie}$ in Equation (71).
@@
Compare to the condition in non-graded geometry by defining the following operations. Let $\rho:=\rho^{i}_{a}(x)\tfrac{\partial}{\partial x^{i}}:E\rightarrow TM$ be a bundle map, $[e_{a},e_{b}]:=C_{ab}^{c}e_{c}$ be a bilinear bracket on $\Gamma(E)$ for the basis $e_{a}\in\Gamma(E)$. Moreover let $J:=\frac{1}{n!}J_{a_{1}\ldots a_{n}}(x)e^{a_{1}}\wedge\ldots\wedge e^{a_{n}}\in\Gamma(\wedge^{n}E^{*})$ be an $E$-$n$-form with the basis $e^{a}\in\Gamma(E^{*})$.
By structures of coordinates, $z_{i}$, $q^{a}$ and $y_{a}$, only nonvanishing terms in ${pr}_{*}e^{\mathrm{ad}\phi}{\widetilde{\Theta}}$ after the projection ${pr}$ are following four terms, $\{\{{\widetilde{\Theta}},{\phi_{Lie}}\},{\phi_{Lie}}\}$, $\{\{{\widetilde{\Theta}},{\underline{J}}{}\},{\phi_{Lie}}\}$, $\{\{{\widetilde{\Theta}},{\phi_{Lie}}\},{\underline{J}}{}\}$ and $\underbrace{\{\ldots\{\{{\widetilde{\Theta}},{\phi_{Lie}}\},{\phi_{Lie}}\},\ldots,{\phi_{Lie}}\}}_{(n+1)\,\text{brackets}}$. At first, $\{\{{\widetilde{\Theta}},{\phi_{Lie}}\},{\phi_{Lie}}\}=0$ is satisfied independently. This equation gives conditions for $\rho^{i}_{a}$ and $C^{a}_{bc}$ that $\rho_{a}=\rho^{i}_{a}\partial_{i}$ is an anchor map and $C^{a}_{bc}$ is a structure function of the Lie algebroid on $E$. Another condition
@@
${pr}_{*}\left(\frac{1}{2}\{\{{\widetilde{\Theta}},{\underline{J}}{}\},{\phi_{Lie}}\}+\frac{1}{2}\{\{{\widetilde{\Theta}},{\phi_{Lie}}\},{\underline{J}}{}\}+\frac{1}{(n+1)!}\underbrace{\{\ldots\{\{\Theta,{\phi_{Lie}}\},{\phi_{Lie}}\},\ldots,{\phi_{Lie}}\}}_{(n+1)\,\text{brackets}}\right)=0,$ (101)
@@
gives the compatible condition of $J$, Equation (11). In fact, the first two terms in (101) is equivalent to ${}^{E}\mathrm{d}J$ and the third term is $\iota_{\rho}^{n+1}H$ by straightforward calculations.
@@
Proposition 7.4
$(E,\rho,[-,-])$ is a Lie algebroid and $J$ is a compatible $E$-$n$-form with a pre-$(n+1)$-form $H$ if and only if $\phi$ satisfies Equation (92) ${pr}_{*}e^{\mathrm{ad}\phi}{\widetilde{\Theta}}=0$, i.e, ${\mathcal{M}}$ is a twisted QP-manifold with $\phi$ in Equations (99) and (100).
@@ Acknowledgments
The author is grateful to the Erwin Schrödinger International Institute for Mathematics and Physics for support within the program “Higher Structures and Field Theory” in 2022, and National Center for Theoretical Sciences and National Tsing Hua University, where part of this work was carried, for their hospitality.
This work was supported by JSPS Grants-in-Aid for Scientific Research Number 22K03323.
He would like to thank to Hsuan-Yi Liao, Camille Laurent-Gengoux and Seokbong Seol for his hospitality and useful discussion. Especially, he would like to thank to to the anonymous referees for relevant contribution to improve the paper.

### v4

@@ 7.2 QP-manifold description of compatible differential forms
Our graded manifold is $T^{*}[n]{\mathcal{M}}=T^{*}[n]T^{*}[n-1]E[1]$. In order to compute the concrete formula, take local coordinates $\bigl{(}x^{i},q^{a},z_{i},y_{a}\bigr{)}$ of degree $(0,1,n-1,n-2)$ on $T^{*}[n-1]E[1]$ and canonical conjugate coordinates of the fiber $\bigl{(}\xi_{i},p_{a},\zeta^{i},\eta^{a}\bigr{)}$ of degree $(n,n-1,1,2)$. The canonical projection ${\operatorname{pr}\colon T^{*}[n]{\mathcal{M}}\rightarrow{\mathcal{M}}}$ is given by $\operatorname{pr}\colon\bigl{(}\xi_{i},p_{a},\zeta^{i},\eta_{a}\bigr{)}\mapsto(0,0,0,0)$. The canonical symplectic form of degree $n$ is
${\widetilde{\omega}}_{\rm grad}$ $=\delta x^{i}\wedge\delta\xi_{i}+\delta q^{a}\wedge\delta p_{a}+\delta z_{i}\wedge\delta\zeta^{i}+\delta y_{a}\wedge\delta\eta^{a}.$
@@
$\big{\{}x^{i},\xi_{j}\big{\}}=-\big{\{}\xi_{j},x^{i}\big{\}}=\delta^{i}_{j},\qquad\{q^{a},p_{b}\}=-(-1)^{n-1}\{p_{b},q^{a}\}=\delta^{a}_{b},$
$\big{\{}z_{i},\zeta^{j}\big{\}}=-(-1)^{n-1}\big{\{}\zeta^{j},z_{i}\big{\}}=\delta_{i}^{j},\qquad\big{\{}y_{a},\eta^{b}\big{\}}=-\big{\{}\eta^{b},y_{a}\big{\}}=\delta_{a}^{b}.$
We take an $(n+1)$-form $H$ on $M$ and consider the following function of degree $n+1$,
@@
$=\xi_{i}\zeta^{i}+p_{a}\eta^{a}+(-1)^{\tfrac{n(n-1)}{2}}\frac{1}{(n+1)!}H_{i_{1}\dots i_{n+1}}(x)\zeta^{i_{1}}\cdots\zeta^{i_{n+1}},$
@@
where $\tilde{H}=\frac{1}{(n+1)!}H_{i_{1}\dots i_{n+1}}(x)\zeta^{i_{1}}\cdots\zeta^{i_{n+1}}$ is a degree $n+1$ function corresponding to the $(n+1)$-form $H=\frac{1}{(n+1)!}H_{i_{1}\dots i_{n+1}}(x)\mathrm{d}x^{i_{1}}\wedge\cdots\wedge\mathrm{d}x^{i_{n+1}}\in\Omega^{n+1}(M)$. An $(n+1)$-form is identified to a degree $n+1$ function on $T[1]M$. ${\widetilde{\Theta}}$ satisfies $\big{\{}{\widetilde{\Theta}},{\widetilde{\Theta}}\big{\}}=0$ if and only if $H$ is closed, $\mathrm{d}H=0$, i.e., $(T^{*}[n]{\mathcal{M}},{\widetilde{\omega}}_{\rm grad},{\widetilde{Q}})$ is a QP-manifold if $H$ is a pre-$n$-plectic form on $M$, where ${\widetilde{Q}}$ is a Hamiltonian vector field for ${\widetilde{\Theta}}$.
@@
We consider a twist using the following function of degree $n$, $\phi={\phi_{\rm Lie}}-{\underline{J}}{}$, where two functions are
${\phi_{\rm Lie}}=\rho^{i}_{a}(x)z_{i}q^{a}+\frac{1}{2}C^{c}_{ab}(x)q^{a}q^{b}y_{c},$ (7.7)
${\underline{J}}{}=\frac{1}{n!}J_{a_{1}\dots a_{n}}(x)q^{a_{1}}\cdots q^{a_{n}}.$ (7.8)
@@
and impose the condition (7.6), $\operatorname{pr}_{*}{\rm e}^{\operatorname{ad}\phi}{\widetilde{\Theta}}=0$. ${\phi_{\rm Lie}}$ is a similar form as $\Theta_{\rm Lie}$ in equation (6.2).
@@
Compare to the condition in non-graded geometry by defining the following operations. Let ${\rho:=\rho^{i}_{a}(x)\tfrac{\partial}{\partial x^{i}}\colon E\rightarrow TM}$ be a bundle map, $[e_{a},e_{b}]:=C_{ab}^{c}e_{c}$ be a bilinear bracket on $\Gamma(E)$ for the basis $e_{a}\in\Gamma(E)$. Moreover let $J:=\frac{1}{n!}J_{a_{1}\dots a_{n}}(x){\rm e}^{a_{1}}\wedge\cdots\wedge{\rm e}^{a_{n}}\in\Gamma(\wedge^{n}E^{*})$ be an $E$-$n$-form with the basis ${\rm e}^{a}\in\Gamma(E^{*})$.
By structures of coordinates, $z_{i}$, $q^{a}$ and $y_{a}$, only nonvanishing terms in $\operatorname{pr}_{*}{\rm e}^{\operatorname{ad}\phi}{\widetilde{\Theta}}$ after the projection $\operatorname{pr}$ are following four terms,
@@
$\big{\{}\big{\{}{\widetilde{\Theta}},{\phi_{\rm Lie}}\big{\}},{\phi_{\rm Lie}}\big{\}},\quad\big{\{}\big{\{}{\widetilde{\Theta}},{\underline{J}}{}\big{\}},{\phi_{\rm Lie}}\big{\}},\quad\big{\{}\big{\{}{\widetilde{\Theta}},{\phi_{\rm Lie}}\big{\}},{\underline{J}}{}\big{\}},\quad\underbrace{\big{\{}\dots\big{\{}\big{\{}{\widetilde{\Theta}},{\phi_{\rm Lie}}\big{\}},{\phi_{\rm Lie}}\big{\}},\dots,{\phi_{\rm Lie}}\big{\}}}_{(n+1)\text{brackets}}.$
At first, $\big{\{}\big{\{}{\widetilde{\Theta}},{\phi_{\rm Lie}}\big{\}},{\phi_{\rm Lie}}\big{\}}=0$ is satisfied independently. This equation gives conditions for $\rho^{i}_{a}$ and $C^{a}_{bc}$ that $\rho_{a}=\rho^{i}_{a}\partial_{i}$ is an anchor map and $C^{a}_{bc}$ is a structure function of the Lie algebroid on $E$. Another condition
$\operatorname{pr}_{*}\bigg{(}\frac{1}{2}\big{\{}\big{\{}{\widetilde{\Theta}},{\underline{J}}{}\big{\}},{\phi_{\rm Lie}}\big{\}}+\frac{1}{2}\big{\{}\big{\{}{\widetilde{\Theta}},{\phi_{\rm Lie}}\big{\}},{\underline{J}}{}\big{\}}$
$\qquad{}+\frac{1}{(n+1)!}\underbrace{\{\dots\{\{\Theta,{\phi_{\rm Lie}}\},{\phi_{\rm Lie}}\},\dots,{\phi_{\rm Lie}}\}}_{(n+1)\,\text{brackets}}\bigg{)}=0$ (7.9)
@@
gives the compatible condition of $J$, equation (3.1). In fact, the first two terms in (7.9) is equivalent to ${}^{E}\mathrm{d}J$ and the third term is $\iota_{\rho}^{n+1}H$ by straightforward calculations.
@@
Proposition 7.4.
$(E,\rho,[-,-])$ is a Lie algebroid and $J$ is a compatible $E$-$n$-form with a pre-$(n+1)$-form $H$ if and only if $\phi$ satisfies equation (7.6) $\operatorname{pr}_{*}{\rm e}^{\operatorname{ad}\phi}{\widetilde{\Theta}}=0$, i.e., ${\mathcal{M}}$ is a twisted QP-manifold with $\phi$ in equations (7.7) and (7.8).
@@ Acknowledgments
The author is grateful to the Erwin Schrödinger International Institute for Mathematics and Physics for support within the program “Higher Structures and Field Theory” in 2022, and National Center for Theoretical Sciences and National Tsing Hua University, where part of this work was carried, for their hospitality. This work was supported by JSPS Grants-in-Aid for Scientific Research Number 22K03323. I would like to thank to Hsuan-Yi Liao, Camille Laurent-Gengoux and Seokbong Seol for his hospitality and useful discussion. Especially, he would like to thank to the anonymous referees for relevant contribution to improve the paper.